A pendulum is fashioned out of a thin bar of length 0.55m and mass 1.9kg, welded to a sphere of radius 0.11m and mass 0.86kg. Find the center of mass of the composite object.
Answer #1

suppose  pendulam  piotal  poin is  origin ( 0,0)

so Center of mass of pendulum = ( 0.55/2 , 0 ) = (0.275 , 0)

and center of mass of sphere from pivot = 0.55+r = 0.55+0.11 = 0.66 m

now center of mass of system

= (m1*x1 +m2*x2) / ( m1+m2)

= ( 1.9*0.275 + 0.86*0.66) /( 1.9+0.86)

=0.394963768 m

so center of mass of the composite object is =0.394963768 m away from pivot

let me know in a comment if there is any problem or doubts